OpenStudy (anonymous):

A radio station has a broadcast area in the shape of a circle with equation x^2 + y^2 = 5,625, where the constant represents square miles. a. Graph the equation and state the radius in miles. b. What is the area of the region in which the broadcast from the station can be picked up?

OpenStudy (raden):

x^2 + y^2 = 5,625 x^2 + y^2 = (75)^5 that is the circle equation with the center at the original point (0,0) and its radius = 75

OpenStudy (anonymous):

oh, I was way off on how I was going to solve this.

OpenStudy (raden):

to get area of that circle, use the formula : area = pi * r * r u just put r=75

OpenStudy (anonymous):

ok, why I think I did that one correct than. Thank you very much.

OpenStudy (raden):

you're welcome

OpenStudy (johnweldon1993):

little typo mistake.... x^2 + y^2 = (75)^5 should be x^2 + y^2 = (75)^2

OpenStudy (johnweldon1993):

doesn't change anything...just a typo mistake

OpenStudy (raden):

thanks johnweldon :)

OpenStudy (johnweldon1993):

no problem :)